# An Empirical Bayes Approach for Estimating Skill Models for Professional Darts Players

### v2

@@ Abstract
We perform an exploratory data analysis on a data-set for the top 16 professional darts players from the 2019 season. We use this data-set to fit player skill models which can then be used in dynamic zero-sum games (ZSGs) that model real-world matches between players. We propose an empirical Bayesian approach based on the Dirichlet-Multinomial (DM) model that overcomes limitations in the data. Specifically we introduce two DM-based skill models where the first model borrows strength from other darts players and the second model borrows strength from other regions of the dartboard. We find these DM-based models outperform simpler benchmark models with respect to Brier and Spherical scores, both of which are proper scoring rules. We also show in ZSGs settings that the difference between DM-based skill models and the simpler benchmark models is practically significant. Finally, we use our DM-model to analyze specific situations that arose in real-world darts matches during the 2019 season.
@@ 1 Introduction
The game of darts has experienced a surge in popularity in recent years and now has a substantial presence in many countries including the U.K., Germany, the Netherlands, and Australia, among others. Indeed, a recent headline in The Economist (2020) explains how darts has moved from pastime to prime time and how it is quite common today to have tens of thousands of fans attend darts tournaments. The article also notes how in recent years darts has become the second-most-watched sport over the Christmas period on Sky Sports in the U.K., coming second only to football. Further growth is expected and professional tournaments now take place in locations such as Shanghai and Madison Square Garden in New York - locations that are far removed from the U.K. and the Netherlands, the traditional hot-beds of darts. Even more recently, in January 2024 the 16-year old Luke Littler became the youngest ever player to reach the final of the PDC World Championships. Beyond the sports pages, his exploits were covered in the New York Times, Bloomberg News and even on the front page of the Financial Times among others. The game is also becoming increasingly attractive to women as evidenced by the exploits of Fallon Sherrock who in 2019 became the first woman to win a match at the PDC World Darts Championship and in fact made it to the third round of that tournament.
@@
The remainder of the paper is organized as follows. In Section 2 we describe the rules of darts and we describe the data-set in Section 3. We propose our DM models in Section 4 and then present the bivariate-normal skill model in Section 5. We use proper scoring rules to compare the DM models versus benchmark models in Section 6 and use dartboard visualizations in Section 7 to emphasize how well the DM models resolve the coverage issues in the data-set. In Section 8 we consider the DM-based skill models in the context of ZSGs and we conclude in Section 9. A Supplementary Material file contains additional material including the EM algorithm for fitting our extension of the bivariate-normal skill model as well as complete versions, i.e. for all 16 players, of all figures and tables that appear in the main text. Finally, we note that some sections, e.g. Sections 2 and 3, closely follow analogous sections in Haugh and Wang (2022).
@@ 2 The Rules of Darts
The rules of darts that we describe here [^3] The rules for 501 and other less common forms of the game are described at https://www.mastersofgames.com/rules/darts-rules.htm. Furthermore the geometry of the dartboard is described in Appendix A of the Supplementary Materials file. are for the most commonly played form of the game, namely “501”. In 501, each player starts on a score of 501 and takes turns in throwing three darts. These three darts constitute a “turn”. After a player’s turn, his scores on the three darts are added together and subtracted from his score at the beginning of the turn. The first player to reach a score of exactly zero wins the game as long as his final dart is a double, i.e. a dart that lands on D1, D2, …, D20, or DB in which case we say the player has checked out. If a player’s turn would result in an updated score of 1, 0 (but not achieved via a double on his final dart) or a negative score, then the turn is invalidated, the player remains on the score he had prior to the turn and his opponent then takes his turn. It’s possible for a player to win without having to throw all three darts in his turn. For example, suppose a player has a score of 16 just prior to his turn. If he then scores D8 with the first dart of his turn he wins. Alternatively, if he scores S8 with his first dart, then he could still win by throwing a D4 with his second dart or (say) S4 and D2 with his second and third darts, respectively.
@@ 3 The Data
We provide the treble data in Appendix E and the entire data-set is available at https://github.com/wangchunsem/OptimalDarts. We now provide some summary statistics of the data. Aggregated across the 16 players, there were 117,600, 27,709, 7,717, and 2,461 attempts at T20, T19, T18, and T17, respectively, and 41.2%, 41.7%, 36.9%, and 33.5% of these attempts, respectively, were successful. However, the success rates at the individual level vary considerably more than the aggregate numbers. For example, several players have a success rate of less than 30% when targeting T18 and over 40% when targeting T20. There were a total of 16,777 attempts at the 21 double regions with D20 (4,399), D16 (2,338), and D10 (2,064) together accounting for over 50% of this total. In contrast, there were only 1,866 attempts in total at the 10 odd doubles. These numbers reflect the preference for even doubles and, in particular, doubles that have powers [^4] If a player needs say D20 to check out on the second dart of his turn, then if he hits an M and or S10 (the most likely fails) he can still check out with his third and final throw in the turn. of 2 in their factorizations. The average success rate at doubles across the 16 players was 40.2%. Note this is smaller than the success rates for T20 and T19, despite the fact that the double beds are approximately 50% larger than the treble areas or “beds”. There are also substantial variations in success rates both across regions for a given player and across players. Some of this variation can certainly be explained by small number effects but even for the most popular double, i.e. D20, we see success rates vary from as low as 27% to as high as 47%.
@@
Unfortunately, the data-set has two important natural limitations. The first is that while we always know the realized value of the score $z$, we never know the precise point $(x,y)$ where each dart lands. This was also true of TPT’s data-set and led to their development of EM algorithms for estimating skill models. The second natural limitation is that when a player targets a particular region, e.g. T20, we do not know the exact point in the region at which he was aiming. This was not true of TPT since they used self-generated data where they aimed at the center of the DB. This second limitation leads to some difficulty in interpreting the skill models from the second stage and in particular, raise the question of bias.
@@ 4 An Empirical Bayes Approach via the DM Distribution (Stage 1)
where ${{\bm{\alpha}}}:=(\alpha_{1},\ldots,\alpha_{K})\in\mathbb{R}^{K}_{+}$ and $\Gamma(\cdot)$ denotes the Gamma function. It is well known that $\text{E}[p_{k}]=\alpha_{k}/\sum_{k}\alpha_{k}$ and that $\text{Var}[p_{k}]=\text{E}[p_{k}](1-\text{E}[p_{k}])/\left(1+\sum_{k}\alpha_{k}\right)$. The Dirichlet distribution is conjugate to the Multinomial distribution which makes the Dirichlet distribution very convenient for modelling proportions.
@@ 4.1 The DM Model: Borrowing Strength from Players
${\bf p}_{j}$ $\sim$ ${\cal D}({{\bm{\alpha}}})$
${\bf z}_{j}$ $\sim$ $\mbox{Mult}(n_{j},{\bf p}_{j})$
@@
Let $D:=\{{\bf z}_{1},\ldots,{\bf z}_{16}\}$ denote the data from the 16 professionals when targeting the region in question. The likelihood is then given by
@@ 4.2 The DM Model: Borrowing Strength from Target Regions
We will describe our approach for the doubles regions and note that the trebles are fitted similarly. Consider, for example, the D20 target which has 7 possible outcomes represented in the data-set: D20, S20, D1, S1, D5, S5, and M (miss). We collapse these 7 outcomes into the following 5 outcomes: D20, S20, {D1,D5}, {S1,S5}, and M. We repeat this for all players and all double targets (D1, …, D20). For each player, we then apply the DM model as in Section 4.1 but now with $K=5$ and $J=20$ regions. The DM model results in new pseudo-counts for each double in the lower $K=5$ dimensional setting. The final step is to split the pseudo-counts (in the case of target D20) for {D1,D5} into pseudo-counts for D1 and D5, and split the pseudo-counts for {S1,S5} into pseudo-counts for S1 and S5. The simplest and probably best way to do this is to divide the new additional counts equally. So for example, if S1 and S5 had 3 and 2 hits in the raw data, respectively, when the target was D20, and we have 3 new pseudo-counts for {S1,S5} then we would divide these 3 equally between S1 and S5 resulting in pseudo-counts of 3+1.5=4.5 and 2+1.5=3.5 for S1 and S5, respectively.
@@ 5 The Bivariate-Normal Skill Model (Stage 2)
In this section we describe the bivariate-normal skill model of TPT and our simple extension of it to accommodate inference of the target ${\bm{\vartheta}}$ (as discussed below). This model can be fitted using either the raw count data or the pseudo-counts from Section 4. The bivariate-normal skill models serves three purposes:
@@
    We require them to fit the benchmark models that we use in Section 6 for evaluating the performance (via proper scoring rules) of the DM models.
@@
    They allow for clear visualisations that enable us to compare the performance of the DM models with models based on the raw data counts. See, for example, many of the figures in Section 7.
@@ 5.1 A Conditional Normal Skill Model
$[x\ y]^{\top}$ $\sim$ $\mbox{N}_{2}({\bm{\mu}}_{i},{{\bm{\Sigma}}}_{i}),\ \ \mbox{ for }{\bm{\mu}}_{i}\in R_{i}$ (5)
@@ 6 Evaluating the DM Models Via Proper Scoring Rules
The first two models are our benchmark models in that neither of them use the empirical Bayes approach of the BSP and BSR models. They are also simpler models in that the Basic model uses the same skill model for the entire dartboard whereas the IJOC model uses the same skill model for all twenty double regions. In contrast, the BSP and BSR models use pseudo-counts to construct a separate skill model for each possible target. In order to fit the Basic and IJOC models we need to apply the EM algorithm (just once for the Basic model and six times for the IJOC model) to estimate the various $({\bm{\mu}},{{\bm{\Sigma}}})$ parameters. As mentioned earlier, our EM algorithm is a simple extension of TPT’s that allows the mean outcome ${\bm{\mu}}$ to also be inferred. Details are in Appendix B. Once the various $({\bm{\mu}},{{\bm{\Sigma}}})$ parameters have been estimated, it is a simple matter of numerical integration to obtain the probabilistic forecasts for use with the scoring rules. In addition to the above four skill models, we also consider two additional skill models:
@@
Models 5 and 6 are therefore fitted using the EM algorithm (with ${\bm{\mu}}$ fixed in Model 6) but applied to the BSP pseudo-counts rather than the raw-data counts [^10] Indeed as noted above it would not be possible to fit a separate skill model using the raw data-counts to each double region due to the scarcity of data for most doubles.. Our main reason for considering the BSPN and BSPNC models is because we may want to be able target very specific locations (down to the square millimeter say) in ZSG settings, while the BSP and BSR skill models do not allow for this. Finally, we note that we could also have considered versions of models 5 and 6 that use the BSR pseudo-counts. We did not consider these models because as we shall see, the BSP and BSR models have very similar performances (in terms of proper scoring rules) with the BSP model appearing to be slightly better than the BSR model.
@@ 6.1 Numerical Results
Before describing our numerical results we first describe how we split the data into training and test sets.
@@
We apply an 80%-20% training-test set split on the entire data-set. Specifically, we aim to take a random sample of 80% of the throws targeted at each player-target region combination as the training-set and take the remaining 20% of throws as the test-set. Because certain targets (D17, D15, D13 and D11) have very few data-points and for some players no data-points at all, we dropped these targets from the analysis. For other rarely player-target combinations where 20% of the number of data-points was less than one, we ensured at least one data-point was in the test-set and then took the remaining points to be in the training-set. We then estimated the various skill models (including the estimation of ${{\bm{\alpha}}}$ for the DM-based models) using the training-set [^11] An alternative, slightly more efficient approach to constructing the train-test set splits would work as follows. We could have used the training-data for player $j$ and all of the data for the remaining players when fitting player $j$’s BSP, BSPN and BSPNC models. We could have done something similar for the BSR model: when evaluating player $j$’s score on a particular target region, e.g. D20, we could have used all of player $j$’s data as the training data except for the throws in the test-set that were aimed at D20. While this would be a more efficient use of the data, it would require more work (e.g. in the number of times ${{\bm{\alpha}}}$ has to be estimated) and so we went with simpler approach outlined above. and estimated the Brier and Spherical scores via the test-set. Finally, since the training-test set splits are random the scores on the test-sets are also random. For this reason we take $N_{s}=20$ random training-test set splits and compute average scores across the $N_{s}$ test sets. Rather than reporting the Brier and Spherical scores for each player-target combination we aggregate them by trebles (T20 through T17), the DB and the 20 other double regions.
@@ Results
We display the average Brier score for all 16 players when targeting the double regions in Table 1. We average across darts rather than regions (D20, D19, etc.) so that each dart in the test set gets equal weight regardless of where it was targeted. Brier score results for the trebles and DB as well as all Spherical score results are reported in Appendix D.1. The Spherical and Brier score results provide very similar rankings of the six skill models which is why we only display the Brier scores here. For each player, we display the highest-ranked score in bold font and do likewise for the average across all players. Our main interest is in the doubles regions since these regions are very important in the game of darts (recall players must check out on a double) but we often only have limited data available for them. Our DM-based estimators were therefore designed with the doubles in mind. (Indeed because the four treble regions have plentiful data the DM-based estimators only provide minimal shrinkage for these regions.)
@@
Our first observation regarding Table 1 is that the BSP, BSR and BSPN models all perform similarly and on average outperform the two benchmark models Basic and IJOC. The BSPNC model also does better than the benchmark models but not as well as the three other DM-based models. This is not surprising since the BSPNC model has two fewer parameters (than the BSPN model) per target region to fit the BSP pseudo-counts. We note that the BSP and BSPN models do marginally better than the BSR models and we also observe this for the Spherical scores reported in Appendix D.1. With regards to the trebles we see that the Basic model performs the worst of all six models and we see that the BSP, BSR and BSPN models are the best performing models with identical (to 4 decimal places) average scores for both Brier and Spherical scores. Given the abundance of data available for the trebles, it is not surprising to see the IJOC model perform well with an average score that is only marginally below the BSP, BSR and BSPN scores. Finally, with regards to the DB we note that there are no results for the BSR model as there are no regions from which the DB region can borrow. For both the Brier and Spherical scores, the BSPN model is ranked highest on the DB followed by the BSPNC and then the Basic model.
@@
We emphasize that the main value of the scoring rules is to provide a ranking of the skill models. In particular, it is difficult to attach much meaning to the differences in reported scores. For example, a proper scoring rule remains a proper scoring rule under any positive affine transformation. Nonetheless, it is possible to provide some interpretation to the numbers in Table 1. For example, when targeting D20 James Wade hits D20, S20, M, D5, S5, D1 and S1 with frequencies of 38.9%, 36.8%, 23.5%, 0%,  0.3%, 0.5%, and 0%, respectively. If we take these frequencies to be the ground truth probability distribution and assume our forecast is perfect, i.e. equal to the ground truth distribution, then the expected Brier score [^12] The reported expected Brier scores in this paragraph are for the case where we have an infinite number of darts. The actual expected Brier score for a finite number of darts will differ due to the non-linearity of the Brier score. will be -0.6580 which would be the best possible Brier score for any skill model for James Wade targeting D20. In contrast, if we used a naive forecast that reported a probability of $1/7$ for each of the 7 possible outcomes, then the resulting expected Brier score would be approximately -0.8571. These values helps to provide some context [^13] James Wade’s results when targeting D20 are typical of the average player in the data-set. That most of the reported numbers for him in Table 1 are slightly better than -0.6580 is due to a number of possible reasons including the fact that the reported numbers are aggregated across all 20 double regions, the non-linearity of the Brier score mentioned in the previous footnote, and random noise due to the variability in the training-test set splits. for the reported Brier scores in Table 1. In Appendix D.2 we argue that the mean Brier and Spherical scores for the DM-based skill models are statistically different from the benchmark models through the use of paired t-tests. Nonetheless, statistical significance does not equate to practical significance. We address practical significance in Section 8.1 where we use our ZSG setting to show that if a player assumes an incorrect skill model (e.g. the Basic or IJOC model) when his true skill model is the BSP model, then his win-probability, i.e. probability of winning a best of $N=35$ leg darts match can decrease by as much as $10\%$ for some players. This follows because a very small difference in forecast quality for a single dart will be magnified and can become practically significant over the many darts thrown in an entire darts match.

### v3

@@ Abstract
We perform an exploratory data analysis on a data-set for the top 16 professional darts players from the 2019 season. We use this data-set to fit player skill models which can then be used in dynamic zero-sum games (ZSGs) that model real-world matches between players. We propose an empirical Bayesian approach based on the Dirichlet-Multinomial (DM) model that overcomes limitations in the data. Specifically we introduce two DM-based skill models where the first model borrows strength from other darts players and the second model borrows strength from other regions of the dartboard. We find these DM-based models outperform simpler benchmark models with respect to Brier and Spherical scores, both of which are proper scoring rules. We also show in ZSGs settings that the difference between DM-based skill models and the simpler benchmark models is practically significant. Finally, we use our DM-based model to analyze specific situations that arose in real-world darts matches during the 2019 season.
@@ 1 Introduction
The game of darts has experienced a surge in popularity in recent years and now has a substantial presence in many countries including the U.K., Germany, the Netherlands, and Australia, among others. Indeed, a recent headline in The Economist (2020) explains how darts has moved from pastime to prime time and how it is quite common today to have tens of thousands of fans attend darts tournaments. The article also notes how in recent years darts has become the second-most-watched sport over the Christmas period on Sky Sports in the U.K., coming second only to football. Further growth is expected and professional tournaments now take place in locations such as Shanghai and Madison Square Garden in New York - locations that are far removed from the U.K. and the Netherlands, the traditional hot-beds of darts. Even more recently, in January 2024 the 16-year old Luke Littler became the youngest-ever player to reach the final of the PDC World Championships. Beyond the sports pages, his exploits were covered in the New York Times, Bloomberg News and even on the front page of the Financial Times among others. The game is also becoming increasingly attractive to women as evidenced by the exploits of Fallon Sherrock who in 2019 became the first woman to win a match at the PDC World Darts Championship and in fact made it to the third round of that tournament.
@@
The remainder of the paper is organized as follows. In Section 2 we describe the rules of darts and we describe the data-set in Section 3. We propose our DM-based models in Section 4 and then present the bivariate-normal skill model in Section 5. We use proper scoring rules to compare the DM-based models versus benchmark models in Section 6 and use dartboard visualizations in Section 7 to emphasize how well the DM-based models resolve the coverage issues in the data-set. In Section 8 we consider the DM-based skill models in the context of ZSGs and we conclude in Section 9. A Supplementary Material file contains additional material including the EM algorithm for fitting our extension of the bivariate-normal skill model as well as complete versions, i.e. for all 16 players, of all figures and tables that appear in the main text. Finally, we note that some sections, e.g. Sections 2 and 3, closely follow analogous sections in Haugh and Wang (2022).
@@ 2 The Rules of Darts
The rules of darts that we describe here [^3] The rules for 501 and other less common forms of the game are described at https://www.mastersofgames.com/rules/darts-rules.htm. Furthermore the geometry of the dartboard is described in Appendix A of the Supplementary Materials file. are for the most commonly played form of the game, namely “501”. In 501, each player starts on a score of 501 and takes turns in throwing three darts. These three darts constitute a “turn”. After a player’s turn, his scores on the three darts are added together and subtracted from his score at the beginning of the turn. The first player to reach a score of exactly zero wins the game as long as his final dart is a double, i.e. a dart that lands on D1, D2, …, D20, or DB in which case we say the player has checked out. If a player’s turn would result in an updated score of 1, 0 (but not achieved via a double on his final dart) or a negative score, then the turn is invalidated, the player remains on the score he had prior to the turn, and his opponent then takes his turn. It’s possible for a player to win without having to throw all three darts in his turn. For example, suppose a player has a score of 16 just prior to his turn. If he then scores D8 with the first dart of his turn he wins. Alternatively, if he scores S8 with his first dart, then he could still win by throwing a D4 with his second dart or (say) S4 and D2 with his second and third darts, respectively.
@@ 3 The Data
We provide the treble data in Appendix E and the entire data-set is available at https://github.com/wangchunsem/OptimalDarts. We now provide some summary statistics of the data. Aggregated across the 16 players, there were 117,600, 27,709, 7,717, and 2,461 attempts at T20, T19, T18, and T17, respectively, and 41.2%, 41.7%, 36.9%, and 33.5% of these attempts, respectively, were successful. However, the success rates at the individual level vary considerably more than the aggregate numbers. For example, several players have a success rate of less than 30% when targeting T18 and over 40% when targeting T20. There were a total of 16,777 attempts at the 21 double regions with D20 (4,399), D16 (2,338), and D10 (2,064) together accounting for over 50% of this total. In contrast, there were only 1,866 attempts in total at the 10 odd doubles. These numbers reflect the preference for even doubles [^4] If a player needs, say D20, to check out on the second dart of his turn, then if he hits an M and or S10 (the most likely fails) he can still check out with his third and final throw in the turn. and, in particular, doubles that have powers of 2 in their factorizations. The average success rate at doubles across the 16 players was 40.2%. Note this is smaller than the success rates for T20 and T19, despite the fact that the double beds are approximately 50% larger than the treble areas or “beds”. There are also substantial variations in success rates both across regions for a given player and across players. Some of this variation can certainly be explained by small number effects but even for the most popular double, i.e. D20, we see success rates vary from as low as 27% to as high as 47%.
@@
Unfortunately, the data-set has two important natural limitations. The first is that while we always know the realized value of the score $z$, we never know the precise point $(x,y)$ where each dart lands. This was also true of TPT’s data-set and led to their development of EM algorithms for estimating skill models. The second natural limitation is that when a player targets a particular region, e.g. T20, we do not know the exact point in the region at which he was aiming. This was not true of TPT since they used self-generated data where they aimed at the center of the DB. This second limitation leads to some difficulty in interpreting the skill models from the second stage and in particular, raises the question of bias.
@@ 4 An Empirical Bayes Approach via the DM Distribution (Stage 1)
where ${{\bm{\alpha}}}:=(\alpha_{1},\ldots,\alpha_{K})\in\mathbb{R}^{K}_{+}$ and $\Gamma(\cdot)$ denotes the Gamma function. It is well known that $\text{E}[p_{k}]=\alpha_{k}/\sum_{k}\alpha_{k}$ and $\text{Var}[p_{k}]=\text{E}[p_{k}](1-\text{E}[p_{k}])/\left(1+\sum_{k}\alpha_{k}\right)$. The Dirichlet distribution is conjugate to the Multinomial distribution which makes the Dirichlet distribution very convenient for modelling proportions.
@@ 4.1 The DM Model: Borrowing Strength from Players
${\bf p}_{j}$ $\sim$ ${\cal D}({{\bm{\alpha}}}),$
${\bf z}_{j}$ $\sim$ $\mbox{Mult}(n_{j},{\bf p}_{j}),$
@@
Let $D:=\{{\bf z}_{1},\ldots,{\bf z}_{16}\}$ denote the data from the 16 professional darts players when targeting the region in question. The likelihood is then given by
@@ 4.2 The DM Model: Borrowing Strength from Target Regions
We will describe our approach for the doubles regions and note that the trebles are fitted similarly. Consider, for example, the D20 target which has 7 possible outcomes represented in the data-set: D20, S20, D1, S1, D5, S5, and M (miss). We combine D20’s double neighbor regions (D1 and D5) and single neighbour regions (S1 and S5), respectively, and hence collapse these 7 outcomes into the following 5 outcomes: D20, S20, {D1,D5}, {S1,S5}, and M. We repeat this for all players and all double targets (D1, …, D20). For each player, we then apply the DM model as in Section 4.1 but now with $K=5$ and $J=20$ regions. The DM model results in new pseudo-counts for each double in the lower $K=5$ dimensional setting. The final step is to split the pseudo-counts (in the case of target D20) for {D1,D5} into pseudo-counts for D1 and D5, and split the pseudo-counts for {S1,S5} into pseudo-counts for S1 and S5. The simplest and probably best way to do this is to divide the new additional counts equally. So for example, if S1 and S5 had 3 and 2 hits in the raw data, respectively, when the target was D20, and we have 3 new pseudo-counts for {S1,S5} then we would divide these 3 equally between S1 and S5 resulting in pseudo-counts of 3+1.5=4.5 and 2+1.5=3.5 for S1 and S5, respectively.
@@ 5 The Bivariate-Normal Skill Model (Stage 2)
In this section we describe the bivariate-normal skill model of TPT and our simple extension of it to accommodate inference of the target ${\bm{\vartheta}}$ (as discussed below). This model can be fitted using either the raw count data or the pseudo-counts from Section 4. The bivariate-normal skill models serve three purposes:
@@
    We require them to fit the benchmark models that we use in Section 6 for evaluating the performance (via proper scoring rules) of the DM-based models.
@@
    They allow for clear visualisations that enable us to compare the performance of the DM-based models with models based on the raw data counts. See, for example, many of the figures in Section 7.
@@ 5.1 A Conditional Normal Skill Model
$[x\ y]^{\top}$ $\sim$ $\mbox{N}_{2}({\bm{\mu}}_{i},{{\bm{\Sigma}}}_{i}),\ \ \mbox{ for }{\bm{\mu}}_{i}\in R_{i},$ (5)
@@ 6 Evaluating the DM Models Via Proper Scoring Rules
The first two models are our benchmark models in that neither of them uses the empirical Bayes approach of the BSP and BSR models. They are also simpler models in that the Basic model uses the same skill model for the entire dartboard whereas the IJOC model uses the same skill model for all twenty double regions. In contrast, the BSP and BSR models use pseudo-counts to construct a separate skill model for each possible target. In order to fit the Basic and IJOC models we need to apply the EM algorithm (just once for the Basic model and six times for the IJOC model) to estimate the various $({\bm{\mu}},{{\bm{\Sigma}}})$ parameters. As mentioned earlier, our EM algorithm is a simple extension of TPT’s that allows the mean outcome ${\bm{\mu}}$ to also be inferred. Details are in Appendix B. Once the various $({\bm{\mu}},{{\bm{\Sigma}}})$ parameters have been estimated, it is a simple matter of numerical integration to obtain the probabilistic forecasts for use with the scoring rules. In addition to the above four skill models, we also consider two additional skill models:
@@
Models 5 and 6 are therefore fitted using the EM algorithm (with ${\bm{\mu}}$ fixed in Model 6) but applied to the BSP pseudo-counts rather than the raw-data counts [^10] Indeed as noted above it would not be possible to fit a separate skill model using the raw data-counts to each double region due to the scarcity of data for most doubles.. Our main reason for considering the BSPN and BSPNC models is that we may want to be able to target very specific locations (down to the square millimeter say) in ZSG settings, while the BSP and BSR skill models do not allow for this. Finally, we note that we could also have considered versions of models 5 and 6 that use the BSR pseudo-counts. We did not consider these models because as we shall see, the BSP and BSR models have very similar performances (in terms of proper scoring rules) with the BSP model appearing to be slightly better than the BSR model.
@@ 6.1 Numerical Results
Before describing our numerical results we first describe how we split the data into training- and test-sets.
@@
We apply an 80%-20% training-test set split on the entire data-set. Specifically, we aim to take a random sample of 80% of the throws targeted at each player-target region combination as the training-set and take the remaining 20% of throws as the test-set. Because certain targets (D17, D15, D13 and D11) have very few data-points and for some players no data-points at all, we dropped these targets from the analysis. For other rarely player-target combinations where 20% of the number of data-points was less than one, we ensured at least one data-point was in the test-set and then took the remaining points to be in the training-set. We then estimated the various skill models (including the estimation of ${{\bm{\alpha}}}$ for the DM-based models) using the training-set [^11] An alternative, slightly more efficient approach to constructing the train-test set splits would work as follows. We could have used the training-data for player $j$ and all of the data for the remaining players when fitting player $j$’s BSP, BSPN and BSPNC models. We could have done something similar for the BSR model: when evaluating player $j$’s score on a particular target region, e.g. D20, we could have used all of player $j$’s data as the training data except for the throws in the test-set that were aimed at D20. While this would be a more efficient use of the data, it would require more work (e.g. in the number of times ${{\bm{\alpha}}}$ has to be estimated) so we went with the simpler approach outlined above. and estimated the Brier and Spherical scores via the test-set. Finally, since the training-test set splits are random the scores on the test-sets are also random. For this reason we take $N_{s}=20$ random training-test set splits and compute average scores across the $N_{s}$ test-sets. Rather than reporting the Brier and Spherical scores for each player-target combination we aggregate them by trebles (T20 through T17), the DB and the 20 other double regions.
@@ Results
We display the average Brier score for all 16 players when targeting the double regions in Table 1. We average across darts rather than regions (D20, D19, etc.) so that each dart in the test-set gets equal weight regardless of where it was targeted. Brier score results for the trebles and DB as well as all Spherical score results are reported in Appendix D.1. The Spherical and Brier score results provide very similar rankings of the six skill models which is why we only display the Brier scores here. For each player, we display the highest-ranked score in bold font and do likewise for the average across all players. Our main interest is in the doubles regions since these regions are very important in the game of darts (recall players must check out on a double) but we often only have limited data available for them. Our DM-based estimators were therefore designed with the doubles in mind. (Indeed because the four treble regions have plentiful data the DM-based estimators only provide minimal shrinkage for these regions.)
@@
Our first observation regarding Table 1 is that the BSP, BSR and BSPN models all perform similarly and on average outperform the two benchmark models Basic and IJOC. The BSPNC model also does better than the benchmark models but not as well as the three other DM-based models. This is not surprising since the BSPNC model has two fewer parameters (than the BSPN model) per target region to fit the BSP pseudo-counts. We note that the BSP and BSPN models do marginally better than the BSR models and we also observe this for the Spherical scores reported in Appendix D.1. With regards to the trebles we see that the Basic model performs the worst of all six models, and we see that the BSP, BSR and BSPN models are the best performing models with identical (to 4 decimal places) average scores for both Brier and Spherical scores. Given the abundance of data available for the trebles, it is not surprising to see the IJOC model perform well with an average score that is only marginally below the BSP, BSR and BSPN scores. Finally, with regards to the DB we note that there are no results for the BSR model as there are no regions from which the DB region can borrow. For both the Brier and Spherical scores, the BSPN model is ranked highest on the DB followed by the BSPNC and then the Basic model.
@@
We emphasize that the main value of the scoring rules is to provide a ranking of the skill models. In particular, it is difficult to attach much meaning to the differences in reported scores. For example, a proper scoring rule remains a proper scoring rule under any positive affine transformation. Nonetheless, it is possible to provide some interpretation to the numbers in Table 1. For example, when targeting D20 James Wade hits D20, S20, M, D5, S5, D1 and S1 with frequencies of 38.9%, 36.8%, 23.5%, 0%,  0.3%, 0.5%, and 0%, respectively. If we take these frequencies to be the ground truth probability distribution and assume our forecast is perfect, i.e. equal to the ground truth distribution, then the expected Brier score [^12] The reported expected Brier scores in this paragraph are for the case where we have an infinite number of darts. The actual expected Brier score for a finite number of darts will differ due to the non-linearity of the Brier score. will be -0.6580 which would be the best possible Brier score for any skill model for James Wade targeting D20. In contrast, if we used a naive forecast that reported a probability of $1/7$ for each of the 7 possible outcomes, then the resulting expected Brier score would be approximately -0.8571. These values help to provide some context [^13] James Wade’s results when targeting D20 are typical of the average player in the data-set. That most of the reported numbers for him in Table 1 are slightly better than -0.6580 is due to a number of possible reasons including the fact that the reported numbers are aggregated across all 20 double regions, the non-linearity of the Brier score mentioned in the previous footnote, and random noise due to the variability in the training-test set splits. for the reported Brier scores in Table 1. In Appendix D.2 we argue that the mean Brier and Spherical scores for the DM-based skill models are statistically different from the benchmark models through the use of paired t-tests. Nonetheless, statistical significance does not equate to practical significance. We address practical significance in Section 8.1 where we use our ZSG setting to show that if a player assumes an incorrect skill model (e.g. the Basic or IJOC model) when his true skill model is the BSP model, then his win-probability, i.e. probability of winning a best of $N=35$ leg darts match can decrease by as much as $10\%$ for some players. This follows because a very small difference in forecast quality for a single dart will be magnified and can become practically significant over the many darts thrown in an entire darts match.